OpenStudy (anonymous):

find equation of plane through the origin and the points <4,-2,5> and <6,1,1>

OpenStudy (calculusfunctions):

Which one. Vector, parametric, symmetric, or scalar equation of the plane?

OpenStudy (anonymous):

simple equation of plane

OpenStudy (experimentx):

ax+by+cz+d=0 --------------- for passing through origin, out x=y=z=0 and get d=0 ---------------------- then put a = 6-4, b=1-(-2) ...

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

find r(t) if r'(t) = 4t i + 3tj +t k and r(1) = i+j

OpenStudy (experimentx):

integrate it ... use the initial value to get rid of constant.

OpenStudy (anonymous):

"then put a = 6-4, b=1-(-2) ... " How does this work?

OpenStudy (anonymous):

If I cross the vectors to get a normal, I get (-7,26,16) so plane is -7(x-4) + 26(y+2) + 16(z-5) = 0 which satisfies all three points given.

OpenStudy (experimentx):

a(x_1) + b(y_1)+ c(z_1) = 0 a(x_2) + b(y_2)+ c(z_2) = 0 ----------------------- a b c ------------ = ----------- = ---------- y1z2 - z1y2 z1x2 - x1z2 x1y2 - y1x2